Manipulate the above equations in the form of y=mx+c and find the slope and y- intercept as well

1.y=aexp(bx)

2.(y+x)(y-x)=a

3.y=acosx^2+bsinx^2

4.yx^2=c


1
Expert's answer
2021-04-14T10:00:15-0400

1.) "y=ae^{bx}"

Taking log on both sides,

"lny=lna+bx\times lne\\lny=lna+bx\\lny=bx+lna"

Let lny=Y and x=X

Y = bX + lna

on comparing we get

Slope = b and y-intercept = lna


2.) "(y+x)(y-x)=a"

"y^2-x^2=a"

Let "y^2=Y and \ x^2=X"

Then, "Y-X=a"

"\Rightarrow Y=X+a"

 On comparing we get

Slope= 1 and y-intercept= a


3.) "y=acos^2x+bsin^2x\\"

"\text{Let cosx=t then}\\y=at^2+b(1-t^2)\\y=(a-b)t^2+b\\y=(a-b)cos^2x+b"

Let y= Y and "cos^2x=X"

Y= (a-b)X +b

On comparing we get

Slope= (a-b) and y-intercept= b


4.)

"yx^2=c\\\text{Taking log on both sides}\\ lny+2lnx=lnc\\ \text{Let lny=Y and lnx=X}\\Y+2X=lnc\\Y=-2X+lnc"

On comparing we get

Slope= -2 and y-intercept= lnc